A $1,000 par value bond was issued five years ago at a 8 percent coupon rate. It currently has 7 years remaining to maturity. Interest rates on similar debt obligations are now 10 percent. Use Appendix B and Appendix D for an approximate answer but calculate your final answer using the formula and financial calculator methods.

A.  Compute the current price of the bond using an assumption of semiannual payments.

B. If Mr. Robinson initially bought the bond at par value, what is his percentage capital gain or loss? [Ignore any interest income receive]

C.  Now assume Mrs. Pinson buys the bond at its current market value and holds it to maturity, what will be her percentage capital gain or loss?
Answer #1

Information given

Par value 1000
Years to maturity 7
Years held till now 5
YTM 10%

a: Using financial calculator
Input: FV= 1000, N=7*2=14, I/Y=10/2=5

PMT=8%*1000/2=40

Solve for PV as -901.01

Current price=$901.01

b: % capital gain/loss =(Selling price-Purchase price)/Purchase price

= (901.01-1000)/1000 = -9.9%

c: % capital gain/loss = (1000-901.01)/901.01 = 10.99%